HUBBLE SPACE TELESCOPE DAILY REPORT #1605 PERIOD COVERED: 0000Z (UTC) 04/03/96 - 0000Z (UTC) 04/04/96 Daily Status Report as of 095/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 6363 (The Environments and Host = Galaxies of BL-Lac Objects) The WF/PC-2 was used to make snapshot survey observations of = BL-Lac object PKS0820+22 in order to study its environment and its host = galaxies. The goal is to investigate systematic trends in host galaxy = type, unusual morphological features, nearby companion galaxies, etc. = The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5999 (The Nuclear Structure of S0 Galaxies: A PC = Imaging Survey) The WF/PC-2 was used to make observations of S0 class galaxy = NGC-1404 in order to study its nuclear structure. This is a part of a = large survey of S0, E/S0, and S0/a class galaxies. The targets will = cover a wide range in luminosity and environment. The observations were = completed as planned, and no anomalies were noted. 1.3 Completed WF/PC-2/FOS(RD)/HRS-1 6736 (A comprehensive Study of H = Ly-Alpha Line Profile and Water) The WF/PC-2, Faint Object Spectrograph (Red Side), and the High = Resolution Spectrograph (Side One) were used to make observations of the = comet Hyakutake in order to make a comprehensive study of hydrogen Lyman = Alpha and water in the comet. The results should improve the = understanding of how water in produced and destroyed in comets. During a = HRS-1 portion of this proposal, the acquisition which occurred at = 094/084023Z failed to single FGS backup on FGS #2. HSTAR 5560 documents = this event (which was reported in yesterday's report). Later, at = 094/162037Z, during another HRS-1 portion, the FGS acquisition failed = when the scan step limit was exceeded. The following acquisition at = 094/175816Z succeeded. HSTAR 5561 covered this event. Otherwise, other = observations in this 17-hour proposal were completed as scheduled, and no = further anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 14 Successful : 13 As documented in HSTAR 5561, the acquisition at 094/162037Z failed = when the scan step limit was exceeded. The following acquisition at = 094/175816Z succeeded. Scheduled Reacquisitions : 2 Successful : 2 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Operations Notes: Operations personnel conducted an on-orbit test for PDB 44.A/PRS = 9.2. Many portions of the test were successful. Some minor problems, = however, were encountered and were documented in HSTAR 5562. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH